What is the equilibrium Co2+ concentration when 1.64 L of a 0.200 M cobalt(II) sulfate solution are mixed with 1.72 L of a 0.357 M potassium carbonate solution?

[Co2+] =____M
Answer #1

Chemical equation between K2CO3 and CoSO4 will be as follows.

CoSO4 + K2CO3 --> CoCO3 + K2SO4

Co2+ + CO32- --> CoCO3

First of all we have to find the number of moles of Co2+ and CO32- available for reaction.

for that we use the equation, Weight = MM'V/1000 where M is the molarity, M' is the molecular mass and V is the volume in ml.

We know Number of moles=(Weight)/(Molecular Mass)

So Number of moles of Co2+ = 0.200*1.64 = 0.328 mole (M=0.200, V=1.64, Molecular mass is canceled)

Number of moles of CO32- = 0.357*1.72 = 0.61404 mole

Then,

(In above picture we can see that,there require 0.328 Co2+ to react with 0.328 CO32- to obtain 0.328 CoCO3 )

​​​​​​​

Thus we can find the equilibrium concentration of Co2+ in the above reaction if we are provided with the fromation consatnt(Kf ) value.

[Co2+] = [CoCO3]/(Kf*[CO32-])

[Co2+] =(0.328)/(0.28604*Kf)

(x in RHS is neglected for convenience)